5) It is known that amounts of money spent on clothing in a year by students on a particular campus follow a normal distribution with a mean of $380 and a standard deviation of $50. a. What is the probability that a randomly chosen student will spend less than $400 on clothing in a year? b. What is the probability that a randomly chosen student will spend more than $360 on clothing in a year? c. What is the probability that a randomly chosen student will spend between $300 and $400 on clothing in a year?


1
Expert's answer
2021-05-31T19:17:29-0400

Solution:

"\mu=380,\sigma=50
\\X\sim N(\mu,\sigma)"

(a) "P(X<400)=P(z<\dfrac{400-380}{50})=P(z<0.4)=0.65542"

(b) "P(X>360)=1-P(X\le360)=1-P(z\le\dfrac{360-380}{50})"

"=1-P(z\le-0.4)=1-[P(z\ge0.4)]=1-[1-P(z\le0.4)]
\\=P(z\le0.4)=0.65542"

(c) "P(300<X<400)=P(X<400)-P(X<300)"

"=P(z<\dfrac{400-380}{50})-P(z<\dfrac{300-380}{50})
\\=P(z<0.4)-P(z<-1.6)
\\=P(z<0.4)-[1-P(z\le1.6)]
\\=0.65542-1+0.94520
\\=0.60062"